D/E Split Night UT 20180925 (first half)

Observers/ISA: B. Rothberg, O. Kuhn (Tucson)
OSA: S. Allanson

ISp: E. Solheid

Mtn Manager: M. Wagner

AO; G. Taylor

UT 02:29 - Pre-opening AO checkout complete, testing LBC+LUCI true Binocular script - IRAF worked already

UT 02:30 - preset accepted successfully - slewing to target after sending scripts separately on each side

UT 02:31 - landed on target (closed dome) - no issues - next to test is when to collimate and send scripts

UT 02:32 - right side failed because guider on right was actually locked onto a hotpixel and tried to move past co-point limit, otherwise all ok so far

UT 03:15 - IT 7399 updated - having issues bringing up blue trackers, have tried various methods, currently warming up electronics before another try

UT 03:25 - trying to turn on LBCs now

UT 03:29 - worked - powered on OK

UT 04:30 - LBC-Blue officially unavailble due to swing arm issue (IT 6891)

UT 04:45 - slight issue with MODS-2 slitmask sievesnap timed out at horizon, retry worked, both MODS now awake.

UT 05:30 - humidity at 99.5% full cloud cover

UT 07:02 - still closed for high humidity and closed - closed GEIRS windows so iNAF can take them.

UT 07:12 - handed over to INAF -- done

-- %USERSIG{BarryRothberg - 2018-09-25}%
Topic revision: r4 - 25 Sep 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback